40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Aug 5 13:28:43 2018, gautam, Update, CDS, c1lsc flaky 
    Reply  Mon Aug 6 00:26:21 2018, gautam, Update, CDS, More CDS woes CDSoverview.png
       Reply  Mon Aug 6 14:38:38 2018, gautam, Update, CDS, More CDS woes CDScrash.pngMC1failures.png
          Reply  Mon Aug 6 19:49:09 2018, gautam, Update, CDS, More CDS woes IMG_7106.JPG
             Reply  Tue Aug 7 11:30:46 2018, gautam, Update, CDS, More CDS woes 
                Reply  Tue Aug 7 22:28:23 2018, gautam, Update, CDS, More CDS woes 
                   Reply  Wed Aug 8 23:03:42 2018, gautam, Update, CDS, c1lsc model started 
                      Reply  Thu Aug 9 12:31:13 2018, gautam, Update, CDS, CDS status update 
                         Reply  Wed Aug 15 21:27:47 2018, gautam, Update, CDS, CDS status update 
                            Reply  Tue Sep 4 10:14:11 2018, gautam, Update, CDS, CDS status update 
                               Reply  Wed Sep 5 10:59:23 2018, wgautam, Update, CDS, CDS status update 
                                  Reply  Thu Sep 6 14:21:26 2018, gautam, Update, CDS, ADC replacement in c1lsc expansion chassis CDSoverview.png
                                     Reply  Fri Sep 7 12:35:14 2018, gautam, Update, CDS, ADC replacement in c1lsc expansion chassis Screenshot_from_2018-09-07_12-34-52.png
                                        Reply  Mon Sep 10 12:44:48 2018, Jon, Update, CDS, ADC replacement in c1lsc expansion chassis 
                                     Reply  Thu Sep 20 11:29:04 2018, gautam, Update, CDS, New PCIe fiber housed PCIeFiberSwap.pngPCIeFiberSwap_FBrebooted.png
                                        Reply  Thu Sep 20 16:19:04 2018, gautam, Update, CDS, New PCIe fiber install postponed to tomorrow 
                                           Reply  Fri Sep 21 16:46:38 2018, gautam, Update, CDS, New PCIe fiber installed and routed PCIeFiber.pngIMG_5878.JPG
Message ID: 14149     Entry time: Thu Aug 9 12:31:13 2018     In reply to: 14146     Reply to this: 14166
Author: gautam 
Type: Update 
Category: CDS 
Subject: CDS status update 

The model seems to have run without issues overnight. Not completely related, but the MC1 shadow sensor signals also don't show any abnormal excursions to negative values in the last 48 hours. I'm thinking about re-connecting the satellite box (but preserving the breakout setup at 1X6 for a while longer) and re-locking the IMC. I'll also start c1ass on the c1lsc frontend. I would say that the other models on c1lsc (i.e. c1oaf, c1cal, c1daf) aren't really necessary for basic IFO operation.

Quote:

As part of this slow but systematic debugging, I am turning on the c1lsc model overnight to see if the model crashes return.

ELOG V3.1.3-